# Linearly-scalable and entropy-optimal learning of nonstationary and nonlinear manifolds

### v1

@@ Abstract
Unsupervised extraction of relevant low-dimensional manifolds from high-dimensional data is in core of many data analysis problems. Common linear methods like the Principal Component Analysis (PCA) and related linear approaches scale linearly with the data statistics size $T$ - but frequently fail to extract the nonlinear or changing in time (nonstationary) low-dimensional manifolds. Common nonlinear methods, like the t-distributed Stochastic Neighbour Embedding (t-SNE), Uniform Manifold Approximation and Projection (UMAP) - as well as their kernelized and parametric extensions (e.g., neural networks for parametric UMAP) - scale as $\mathcal{O}(T^{2})$, or $\mathcal{O}(T\textrm{log}(T))$ in the best case, frequently fail extracting nonstationary manifolds, and can distort results by generating method-induced artefacts (e.g., reveal clusters that are actually not present in the data, or do not reveal clusters when they are present, some examples are provided). PCA-based clustering methods scale as $\mathcal{O}(T)$ in every iteration, and can be used to extract linear nonstationary manifolds, but fail in the situations when the manifolds are nonlinear (some examples are also provided). Here we propose an Entropy-Optimal Manifold Clustering (EOMC) as a metricised and entropy-regularized extension of PCA clustering - and show that it mitigates the problems of the existing tools even in very nonstationary and nonlinear situations, while pertaining the favourable $\mathcal{O}(T)$ iteration complexity scaling. In comparison with the state-of-the-art linear and nonlinear methods on a set of noisy high-dimensional synthetic benchmarks, EOMC is demonstrated to provide a very robust artefact-free learning and reconstruction of low-dimensional manifolds from noisy, nonlinear and nonstationary data. Application to the Lorenz-96 dynamical system - a very popular model of a turbulent behaviour in one dimension - in chaotic and strongly-chaotic regimes reveals that its dynamics is essentially described by a metastable regime-switching process, making infrequent transitions between the very persistent three-dimensional attractive manifolds. The dimensionality of these manifolds appears to remain unchanged, and their overall number gradually grows with the growing external forcing of the Lorenz-96 model. At the same time, the Markovian mean exit times and relaxation times (that bound the predictability horizons for the identified regime-switching process) appear to decrease only very slowly with the growing external forcing - indicating approximately two-fold longer prediction horizons then is currently anticipated based on analysis of positive Lyapunov exponents for this system, even in very chaotic model regimes. It is also demonstrated that when applied for a lossy compression of the Lorenz-96 output data in various forcing regimes, EOMC achieves several orders of magnitude smaller compression loss - when compared to the common PCA-related linear compression approaches that build a backbone of the state-of-the-art lossy data compression tools (like JPEG, MP3, and others). These findings open new exciting opportunities for EOMC and transfer operator theory, by improving predictive skills and performance of data-driven tools in fluid mechanics and geosciences applications.
@@ 2.1 Mathematical formulation of the PCA-clustering
From Lemmas 1 and 2 follows the first major problem of PCA-clustering: according to the Lemmas, the kernel of the projection operator is non-empty when $d<D$ (Lemma 1), and any perturbations inside of the kernel do not have any effect on the loss (Lemma 2). Hence, loss function in (2.1) is a semi-norm, not penalizing errors in the directions that are orthogonal, or close to orthogonal with respect to the manifolds $\mathcal{T}_{k}$. As will be demonstrated on examples below (see, e.g., Fig. 2C and the left panel of Fig. 3C), this can lead to the PCA-clustering minimizers $\mathcal{T}_{k}$ that become orthogonal, or close to orthogonal, with respect to the actual nonlinear low-dimensional manifold - hereby ignoring the actual manifold data, and achieving a small value of $\mathcal{L}$ at the same time. The second problem of PCA-clustering is induced by the exact $0/1$ minimizer in the step (ii) described above: it confines the resulting manifold reconstruction $X^{\textrm{rec}}(:,t)$ in (2) to piecewise-linear functions only, not allowing to achieve smooth convex interpolations of nonlinear manifolds that should be possible with the original optimization problem formulation (2.1). But, this original formulation (2.1) does not allow obtaining the cheaply-computable analytic linearly-scalable solutions (i)-(ii) that are used in the PCA-clustering algorithm. And, solving the original problem (2.1) - being highly-nonlinear, subject to multivariate and polynomial constraints - with standard numerical tools from optimization theory would require very expensive numerical algorithms that have the exponential worst-case scaling [19].
@@ 2.2 Entropy-Optimal Manifold Clustering (EOMC)
  • Entropy regularization: to avoid the problem with a piecewise-linearity restriction - and to allow for the $\gamma$-solutions to be not only $0/1$ (and, to achieve this without returning back to the original formulation (2.1) that would result in exponentialy-scaling worst case numerics) - we can use the fact that the columns of $\gamma$ can be interpreted as $t$-dependent $K$-dimensional probability measures. We can add the normalized negative Shannon entropy regularization terms $\beta\gamma(k,t)\textrm{log}_{K}\left(\gamma(k,t)\right)$ (with a scalar hyper-parameter $\beta\geq 0$) to the loss, to optimally-adjust the entropies of these probability measures during learning [^2] Base $K$ of the logarithm is used to normilize the entropy values to the interval $\left[0,1\right]$. This modification is motivated by the entropic learning methods that allow training the entropy-optimal probability measures by tuning this hyper-parameter $\beta$ in the loss. The $0/1$ $\gamma$-solution of the PCA-clustering is a minimum entropy measure, and is attained when setting $\beta=0$ [20, 21, 22, 23, 24].
@@
2.3 Selection of EOMC hyper-parameters $d,K,\alpha,\beta$
@@
In comparison with the state-of-the-art nonlinear manifold learning methods like t-SNE and UMAP, EOMC relies on a much smaller set of hyper-parameters. It requires to tune only the reduced manifold dimensionality $d$, the number $K$ of locally-linear manifolds for nonlinear approximation, as well as two non-negative scalar regularization parameters $\alpha$ and $\beta$. In contrast, methods like t-SNE require tuning a much larger set of hyper-parameters, that besides of the reduced manifold dimensionality parameter $d$ and number of PCA dimensions $K$, include such model-specific adjustable parameters like perplexity, exaggeration, and Barnes-Hut tradeoff parameter. In addition, UMAP model-specific hyper-parameters include the number of neighbours, metric and its weight, and minimal distance between embedded points. Moreover, numerics of these nonlinear manifold learning methods relies on (stochastic) gradient descent algorithm - that requires tuning of multiple additional hyper-parameters that can have a very strong effect on convergence and cost of the learning phase (like learning rate schedule, batch size, cache size, etc.). As shown in the Theorem 1 in the previous section, EOMC is performed without deploying the (stochastic) gradient numerics. Instead, EOMC performs an iterative repetition of two analytic solutions in Step 1 and Step 2, not requiring additional hyper-parameter tuning.
@@
As will be demonstrated below on practical examples (for example, see Fig. 1), tuning the EOMC hyper-parameters $d$ and $K$ can be done iteratively. To select the optimal values for $d$ and $K$, for each of the $k=1,\dots,K$ manifolds one inspects the decay of eignevalues $\Lambda^{(k)}_{1,1},\Lambda^{(k)}_{2,2},\dots,\Lambda^{(k)}_{D,D}$ for the local weighted covariances $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ in the local manifolds: as follows from the Theorem 1 above, $\sum_{i=d+1}^{D}\Lambda^{(k)}_{i,i}$ quantifies the amount of squared 2-norm loss from projecting the data, that is "assigned" to this manifold $k$, through the EOMC internal coordinates $\gamma(k,:)$. Hence, it quantifies the local quality of lossy compression, that is achieved when projecting the high-dimensional data on the local low-dimensional manifolds. Since (2.2) is an unsupervised learning problem, we can not apply most of the standard hyper-paremeter selection routines from ML and AI, like cross-validation and Bayesian hyper-paremeter tuning [25]. Instead, one can use here the hyper-parameter selection from the unsupervised regularization methods in statistics and computational science, like the $L$-curve method [26]. Alternatively, in the following examples we will adopt a nonparametric information-theoretic perspective to model selection - aiming to find the hyper-parameter combinations that lead to the models combining simplicity (measured as a high lossy compression rate, computed as a ratio between the raw data complexity and EOMC model descriptor length), and quality (measured as the relative loss of compression). In contrast to common parametric measures from information theory (like Akaike and Bayesian Information Criteria) [27] that rely on validity of parametric assumptions like Gaussianity, this non-parametric model selection procedure based on lossy compression allows a robust identification of good hyper-parameter combinations across all of the benchmarks considered below. As will also be shown on application examples below, EOMC does not require a careful and precise adjustment, since the results remain robust in the broad ranges of hyper-parameters.
@@ 3 Application examples
First, we will demonstrate several applications of the EOMC algorithm introduced above, in comparison with the state-of-the-art methods, for the synthetic noisy data examples with known low-dimensional manifold structures, and for the data produced by the Lorenz-96 model from fluid mechanics.
@@ 3.1 Synthetic examples
3.2 Example 1: nonstationary mixture of data switching between 3D ball surface and 3D torus surface in 10 dimensions with noise
Figure 1: Graphic illustration of the five phases of the EOMC data analysis pipeline. Text description is provided in the Sec. 3.2.
@@
  • Phase 1: select the hyper-parameters $d,K,\alpha,\beta$ as described in Sec. 2.3 and perform the EOMC analysis according to the Theorem 1 above. Note that the application of the standard linear PCA (left panel in Fig. 1) does not reveal any spectral gap in the eigenvalues of the covariance, and does not indicate a presence of any low-dimensional manifold. In contrast, applying EOMC with arbitrarily selected $K=15$ indicates a clear spectral gap after the second eigenvalues of the localized EOMC covariance matrices $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$. This clearly indicates that the data contains non-linear manifolds with local dimensionality $d=2$.
@@
3.3 Example 2: nonstationary mixture of data switching between 1D peace sign contour (planar) and 1D prism contour (in 3D), embedded and rotated in 100 dimensions with noise
@@
Next, we will consider a non-stationary process switching between the two non-linear one-dimensional manifolds: a 1D contour of a 3D prism and a 1D contour of the 2D peace sign, containing both piece-wise linear and nonlinear parts. To make the problem even more challenging for manifold learning (then in the previous example from Sec. 3.2), generated 3D data matrix is further randomly rotated in 100 dimensions and subject to a 100-dimensional Gaussian noise.

### v2

@@ Abstract
Unsupervised extraction of relevant low-dimensional manifolds from high-dimensional data is in core of many data analysis problems. Common linear methods like the Principal Component Analysis (PCA) and related linear approaches scale linearly with the data statistics size $T$ - but frequently fail to extract the nonlinear or changing in time (nonstationary) low-dimensional manifolds. Common nonlinear methods scale as $\mathcal{O}(T^{2})$, or $\mathcal{O}(T\textrm{log}(T))$ in the best case, frequently fail extracting nonstationary manifolds, and can distort results by generating method-induced artefacts (e.g., reveal clusters that are actually not present in the data, or do not reveal clusters when they are present, some examples are provided). Here we propose an Entropy-Optimal Manifold Clustering (EOMC) as a metricised and entropy-regularized extension of PCA clustering - and show that it mitigates the problems of the existing tools even in very nonstationary and nonlinear situations, while pertaining the favourable $\mathcal{O}(T)$ iteration complexity scaling and allowing cheap explicit computation of data reliability, i.e., the probability that the data comes from a definition domain of the learned manifold mapping. In comparison with the state-of-the-art linear and nonlinear methods on a set of noisy high-dimensional synthetic benchmarks, EOMC is demonstrated to provide a remarkably robust artefact-free learning and reconstruction of low-dimensional manifolds from noisy, nonlinear and nonstationary data. Application to the Lorenz-96 dynamical system in chaotic regime, as well as to a modified Hasegawa-Wakatani (mHW) model of drift-wave turbulence in the edge of a tokamak plasma reveals that for both of the models their essential dynamics is best described as a metastable regime-switching process, making infrequent transitions between the very persistent low-dimensional manifolds. At the same time, the Markovian mean exit times and relaxation times (that bound the predictability horizons for the identified regime-switching process) appear to decrease only very slowly with the growing external forcing - indicating approximately two-fold longer prediction horizons then is currently anticipated based on analysis of positive Lyapunov exponents, even in very chaotic model regimes. It is also demonstrated that when applied for a lossy compression of the Lorenz-96 and mHW output data in various forcing regimes, EOMC achieves several orders of magnitude smaller compression loss - when compared to the common PCA-related linear compression approaches that build a backbone of the state-of-the-art lossy data compression tools (like JPEG, MP3, and others). These findings open new exciting opportunities for EOMC and transfer operator theory, by offering new possibilities to significantly improve predictive skills and performance of data-driven tools in fluid mechanics and geosciences applications.
@@ 2.1 Mathematical formulation of the PCA-clustering
From Lemmas 1 and 2 follows the first major problem of PCA-clustering: according to the Lemmas, the kernel of the projection operator is non-empty when $d<D$ (Lemma 1), and any perturbations inside of the kernel do not have any effect on the loss (Lemma 2). Hence, loss function in (2.1) is a semi-norm, not penalizing errors in the directions that are orthogonal, or close to orthogonal with respect to the manifolds $\mathcal{T}_{k}$. As will be demonstrated on examples below (see, e.g., Fig. 2C and the left panel of Fig. 3C), this can lead to the PCA-clustering minimizers $\mathcal{T}_{k}$ that become orthogonal, or close to orthogonal, with respect to the actual nonlinear low-dimensional manifold - hereby ignoring the actual manifold data, and achieving a small value of $\mathcal{L}$ at the same time. The second problem of PCA-clustering is induced by the exact $0/1$ minimizer in the step (ii) described above: it confines the resulting manifold reconstruction $X^{\textrm{rec}}(:,t)$ in (2) to piece-wise linear functions only, not allowing to achieve smooth convex interpolations of nonlinear manifolds that should be possible with the original optimization problem formulation (2.1). But, this original formulation (2.1) does not allow obtaining the cheaply-computable analytic linearly-scalable solutions (i)-(ii) that are used in the PCA-clustering algorithm. And, solving the original problem (2.1) - being highly-nonlinear, subject to multivariate and polynomial constraints - with standard numerical tools from optimization theory would require very expensive numerical algorithms that have the exponential worst-case scaling [19].
@@ 2.2 Entropy-Optimal Manifold Clustering (EOMC)
  • Entropy regularization: to avoid the problem with a piece-wise linearity restriction - and to allow for the $\gamma$-solutions to be not only $0/1$ (and, to achieve this without returning back to the original formulation (2.1) that would result in exponentialy-scaling worst case numerics) - we can use the fact that the columns of $\gamma$ can be interpreted as $t$-dependent $K$-dimensional probability measures. We can add the normalized negative Shannon entropy regularization terms $\beta\gamma(k,t)\textrm{log}_{K}\left(\gamma(k,t)\right)$ (with a scalar hyper-parameter $\beta\geq 0$) to the loss, to optimally-adjust the entropies of these probability measures during learning [^2] Base $K$ of the logarithm is used to normilize the entropy values to the interval $\left[0,1\right]$. This modification is motivated by the entropic learning methods that allow training the entropy-optimal probability measures by tuning this hyper-parameter $\beta$ in the loss. The $0/1$ $\gamma$-solution of the PCA-clustering is a minimum entropy measure, and is attained when setting $\beta=0$ [20, 21, 22, 23, 24].
@@
2.3 Incorporating simultaneous learning of data reliability in EOMC
Learning piece-wise low-dimensional linear manifold approximations with EOMC implicitly deploys an assumption that all of the data instances have the same uniform probability $(1/T)$ to contribute to the loss function $\mathcal{L}^{\textrm{EOMC}}$ in (2.2). This is not a problem when this assumption is justified, and when the training data is spread more-or-less uniformly and close to the manifold. But, this becomes a problem when the new data instances $X$ (f.e., the ones that were not used in the training), come from somewhere arbitrarily in the feature space - and further away from the manifold. In the Sec.3.1.3 below, we provide an example of such a situation - resulting in projection artefacts, when training the manifold with the data close to the manifold (red crosses in Fig. 4), and then using the trained EOMC manifold projectors (11) from Theorem 1 for the data points that are far away from the manifold and from the training data (blue dots in Fig. 4B). This input data reliability issue is a very general and severe problem in contemporary AI, resulting in low robustness, "over-confidence" in provided answers, and in AI vulnerability to the so-called adversarial attacks, when subject to new input data that is sufficiently-different from the domain $\Omega$ of "admissible" data implicitly defined by the training data.
To address this issue - and following the ideas originally introduced in the Entropic Outlier Sparsification (EOS) and Entropy-Optimal Networks (EON) methods [21, 24] - we propose to incorporate in the EOMC problem formulation (2.2) an additional learnable variable $\gamma_{0}(t)=\mathbb{P}\left[X(:,t)\in\Omega\right]$, measuring a probability that a new data instance $X(:,t)$ belongs to the "admissible" data domain $\Omega$. Denoting $l(t)=-\sum_{k=1}^{K}\gamma(k,t)\left[\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)+\beta\textrm{log}_{K}\left(\gamma(k,t)\right)\right]$, we can now rewrite (2.2) as a maximisation of the expectation $\mathbb{E}_{\gamma_{0}}\left[l\right]$ of the loss $l$ wrt. this probability measure $\gamma_{0}$, subject to the Shannon entropy regularization terms $\beta_{0}\gamma_{0}\left(t\right)\textrm{log}_{T}\gamma_{0}\left(t\right)$, where $\beta_{0}\geq 0$. Re-arranging terms we obtain:
$\left\{\gamma_{0}^{*},\gamma^{*},\mu^{*}_{1},\mathcal{T}^{*}_{1},\dots,\mu^{*}_{K},\mathcal{T}^{*}_{K}\right\}=\arg\min\mathcal{L}_{\gamma_{0}}^{\textrm{EOMC}}$
$\mathcal{L}_{\gamma_{0}}^{\textrm{EOMC}}$ $=$ $\sum_{t=1}^{T}\gamma_{0}(t)\left[\beta_{0}\textrm{log}_{T}\gamma_{0}\left(t\right)-l(t)\right],$
s.t. $l(t)=-\sum_{k=1}^{K}\gamma(k,t)\left[\left(X(:,t)-\mu_{k}\right)^{\dagger}\left(\left(I+\alpha\right)-\mathcal{T}_{k}\mathcal{T}_{k}^{\dagger}\right)\left(X(:,t)-\mu_{k}\right)+\beta\textrm{log}_{K}\left(\gamma(k,t)\right)\right],$ (14)
$\mathcal{T}_{k}^{\dagger}\mathcal{T}_{k}=I_{d},$
$\gamma(k,t)\geq 0,\quad\textrm{and }\sum_{k=1}^{K}\gamma(k,t)=1,\quad\forall t,k,$
$\gamma_{0}(t)\geq 0,\quad\textrm{and }\sum_{t=1}^{T}\gamma_{0}(t)=1,\quad\forall t.$
Shannon entropy regularization terms $\beta_{0}\gamma_{0}\left(t\right)\textrm{log}_{T}\gamma_{0}\left(t\right)$ are incorporated here with a positive sign, since for $\beta_{0}\equiv 0$, $\gamma_{0}$ solution of (2.3) is entropy-minimal, attaining only $1/0$ values - and the entropy-optimal $\gamma_{0}$-solutions can only be found when maximizing the Shannon entropy. It is also straightforward to validate that the previous formulation (2.2) becomes a special case of (2.2) when $\beta_{0}\rightarrow+\infty$ (since then $\gamma_{0}(t)\rightarrow(1/T)$, for all $t$).
Then, following the same line of argumentation as in the proof of Theorem 1 above, it is easy to show that for the fixed values of the loss $l$ (hence, when all other learning variables except of $\gamma_{0}$ are fixed), and for any $\beta_{0}>0$, there exists a unique solution $\gamma^{*}_{0}$ of the problem (2.3):
$\gamma_{0}^{*}(t)$ $=$ $\frac{\exp\left(\beta_{0}^{-1}\left(l_{\max}+l(t)\right)\right)}{\sum_{t=1}^{T}\exp\left(\beta_{0}^{-1}\left(l_{\max}+l(t)\right)\right)},$ (15)
where $l_{\max}(t)$ is the maximum of $l$ over all $t$. This solution can be computed with an additional cost of $\mathcal{O}(T)$ - independent of $K$ and $D$ (since the values of loss $l$ are pre-computed in the original EOMC algorithm and can be recycled). Hence, adding Step 3 (15) to the original algorithm described in the Theorem 1 above can be done essentially for free, not increasing the leading order of the EOMC algorithm complexity, and not violating the monotonicity of EOMC convergence.
2.4 Selection of EOMC hyper-parameters $d,K,\alpha,\beta,\beta_{0}$
@@
In comparison with the state-of-the-art nonlinear manifold learning methods like t-SNE and UMAP, EOMC relies on a smaller set of hyper-parameters. It requires to tune only the reduced manifold dimensionality $d$, the number $K$ of locally-linear manifolds for nonlinear approximation, as well as two or three non-negative scalar regularization parameters $\alpha$, $\beta$ (and $\beta_{0}$, if simultaneous learning of data reliability is incorporated in EOMC, as described in Sec. 2.3). In contrast, methods like t-SNE require tuning a much larger set of hyper-parameters, that besides of the reduced manifold dimensionality parameter $d$ and number of PCA dimensions $K$, include such model-specific adjustable parameters like perplexity, exaggeration, and Barnes-Hut tradeoff parameter. In addition, UMAP model-specific hyper-parameters include the number of neighbours, metric and its weight, and minimal distance between embedded points. Moreover, numerics of these nonlinear manifold learning methods relies on (stochastic) gradient descent algorithm - that requires tuning of multiple additional hyper-parameters that can have a very strong effect on convergence and cost of the learning phase (like learning rate schedule, batch size, cache size, etc.). As shown in the Theorem 1 above, EOMC is performed without deploying the (stochastic) gradient numerics. Instead, EOMC performs an iterative repetition of analytic solutions not requiring additional numerics-specific hyper-parameter tuning in Step 1, Step 2 (and in additional Step 3 defined by (15), if data reliability is also learned).
@@
As will be demonstrated below on practical examples (for example, see Fig. 1), tuning the EOMC hyper-parameters $d$ and $K$ can be done iteratively. To select the optimal values for $d$ and $K$, for each of the $k=1,\dots,K$ manifolds one inspects the decay of eigenvalues $\Lambda^{(k)}_{1,1},\Lambda^{(k)}_{2,2},\dots,\Lambda^{(k)}_{D,D}$ for the local weighted covariances $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$ in the local manifolds: as follows from the Theorem 1 above, $\sum_{i=d+1}^{D}\Lambda^{(k)}_{i,i}$ quantifies the amount of squared 2-norm loss from projecting the data, that is "assigned" to this manifold $k$, through the EOMC internal coordinates $\gamma(k,:)$. Hence, it quantifies the local quality of lossy compression, that is achieved when projecting the high-dimensional data on the local low-dimensional manifolds. Since (2.2) is an unsupervised learning problem, we can not apply most of the standard hyper-paremeter selection routines from ML and AI, like cross-validation and Bayesian hyper-paremeter tuning [25]. Instead, one can use here the hyper-parameter selection from the unsupervised regularization methods in statistics and computational science, like the $L$-curve method [26]. Alternatively, in the following examples we will adopt a nonparametric information-theoretic perspective to model selection - aiming to find the hyper-parameter combinations that lead to the models combining simplicity (measured as a high lossy compression rate, computed as a ratio between the raw data complexity and EOMC model descriptor length), and quality (measured as the relative loss of compression). In contrast to common parametric measures from information theory (like Akaike and Bayesian Information Criteria) [27] that rely on validity of parametric assumptions like Gaussianity, this non-parametric model selection procedure based on lossy compression allows a robust identification of good hyper-parameter combinations across all of the benchmarks considered below. As will also be shown on application examples below, EOMC does not require a careful and precise adjustment, since the results remain robust in the broad ranges of hyper-parameters.
@@ 3 Application examples
Next, we will demonstrate several applications of the EOMC algorithm introduced above, in comparison with the state-of-the-art methods, for the synthetic noisy data examples with known low-dimensional manifold structures, for the data produced by the simplified Lorenz-96 model from fluid mechanics, as well as for the much more realistic modified Hasegawa-Wakatani (mHW) model of drift-wave turbulence in the edge of a tokamak plasma.
@@ 3.1 Synthetic examples
3.1.1 Example 1: nonstationary mixture of data switching between 3D ball surface and 3D torus surface in 10 dimensions with noise
Figure 1: Graphic illustration of the five phases of the EOMC data analysis pipeline. Text description is provided in the Sec. 3.1.1.
@@
  • Phase 1: select the hyper-parameters $d,K,\alpha,\beta$ (and $\beta_{0}$, if simultaneous learning of data reliability $\gamma_{0}$ with (2.3) is incorporated in EOMC [^3] Here, and in the next example described in Sec. 3.1.2, we perform the analysis setting $\gamma_{0}\equiv 1/T$ and $\beta_{0}=0$. An example of EOMC with simultaneous learning of the entropy-optimal data reliability function $\gamma_{0}$ will be provided in Sec. 3.1.3), as described in Sec. 2.4. Next, perform the EOMC analysis according to the Theorem 1 above. Note that the application of the standard linear PCA (left panel in Fig. 1) does not reveal any spectral gap in the eigenvalues of the covariance, and does not indicate a presence of any low-dimensional manifold. In contrast, applying EOMC with arbitrarily selected $K=15$ indicates a clear spectral gap after the second eigenvalues of the localized EOMC covariance matrices $\mathrm{Cov}_{k}\left(X,\gamma,\mu_{k}^{*}\right)$. This clearly indicates that the data contains non-linear manifolds with local dimensionality $d=2$.
@@
3.1.2 Example 2: nonstationary mixture of data switching between 1D peace sign contour (planar) and 1D prism contour (in 3D), embedded and rotated in 100 dimensions with noise
@@
Next, we will consider a non-stationary process switching between the two non-linear one-dimensional manifolds: a 1D contour of a 3D prism and a 1D contour of the 2D peace sign, containing both piece-wise linear and nonlinear parts. To make the problem even more challenging for manifold learning (then in the previous example from Sec. 3.1.1), generated 3D data matrix is further randomly rotated in 100 dimensions and subject to a 100-dimensional Gaussian noise.
@@
3.1.3 Example 3: co-inference of input reliability $\gamma_{0}(t)=\mathbb{P}\left[X(:,t)\in\Omega\right]$ in EOMC learning
Next, we illustrate the application of the modified EOMC (with a simultaneous learning the input data reliability $\gamma_{0}$), as described in Sec.  2.3 and eqs. (2.3-15). We generate a noisy 2D data (red crosses in Fig. 4A), that are normally distributed around a smooth 1D non-linear manifold following a smooth letter "S" (black line in Fig. 4A). As in the previous synthetic examples, to make the manifold learning more challenging, we add an additional 10-dimensional Gaussian noise and randomly rotate the training data in 10 dimensions. As discussed in Sec.  2.3, the EOMC formulation in (2.2) implicitly assumes that all of the training data points have the same $1/T$ contribution to the loss function - i.e., according to the uniform probability distribution prior. However, this assumption can be seriously violated when considering the data points not used in training and distributed uniformly inside of the whole 10-dimensional hypercube - and far away from the training data (blue dots in Fig. 4). As can be seen from Fig. 4B, applying the projection formula (11) from the Theorem 1 results in artefact projections that are far away of the true manifold (red circular dots in Fig. 4B).
In contrast, co-inference of data reliability function $\gamma_{0}(t)=\mathbb{P}\left[X(:,t)\in\Omega\right]$ in the modified EOMC from Sec.  2.3 and eqs. (2.3-15) produces a very reliable reconstruction of the true manifold: the intensity of red for circular dots in Fig. 4C is proportional to the inferred values of data reliability $\gamma_{0}(t)$. The unreliable points far away from the training data and from the EOMC-approximated manifold automatically get values of $\gamma_{0}(t)$ close to zero. As in the previous two synthetic examples, achieving this result did not require any particular hyper-parameter tuning: it was obtained for an ad hoc choice of $K=6,d=1,\alpha=5,\beta=0.3,\beta_{0}=4$. This result remains remarkably-robust, and does not change notably when changing these hyper-parameter values in quite broad ranges.
Figure 4: Illustration of the modified EOMC learning from Sec.  2.3 and eqs. (2.3-15): co-inference of data reliability function $\gamma_{0}$.